# Sensitivity of the Global 21-cm Signal to Dark Matter-Baryon Scattering

### v1

@@ II Global 21-cm Signal with IDM
Table 1: Comparison of four experimental setups, outlining their key parameters such as minimum and maximum frequencies probed ($\nu_{\mathrm{min}},\nu_{\mathrm{max}}$), number of independent frequency channels ($\mathrm{N}_{\mathrm{ch}}$), native (not smoothed) frequency bins ($\Delta\nu$), observation times ($\tau_{\mathrm{obs}}$), receiver temperatures ($\mathrm{T}_{\mathrm{rec}}$), and RMS value of the smoothed noise ($\sigma_{\mathrm{RMS}}$).
III Methods
III.1 Fisher Forecasting
$F_{\mathrm{ij}}=-\left<\frac{\partial^{2}\ln\mathcal{L}}{\partial\theta_{\mathrm{i}}\partial\theta_{\mathrm{j}}}\right>\,,$ (8)
where $F_{\mathrm{ij}}$ are the $(\mathrm{i},\mathrm{j})$-th elements of the Fisher matrix, $\theta_{\mathrm{i}}$ are the model parameters, $\ln\mathcal{L}$ represents the log-likelihood function, and <>absent<>< > denotes the ensemble average over different data realizations.
$\mathrm{C}_{\mathrm{ij}}=\mathrm{F}^{-1}_{\mathrm{ij}}\,,$ (9)
where $\mathrm{C}_{\mathrm{ij}}$ are the $(\mathrm{i},\mathrm{j})$-th elements of the covariance matrix and $\mathrm{F}^{-1}_{\mathrm{ij}}$ are the $(\mathrm{i},\mathrm{j})$-th element of the inverse of Fisher matrix. Once the Fisher matrix is calculated, one can compute the uncertainties (standard deviations) of individual parameters $\theta_{\mathrm{i}}$ as follows:
$\sigma(\theta_{\mathrm{i}})=\sqrt{\mathrm{F}^{-1}_{\mathrm{ii}}}\,,$ (10)
$\rho_{\mathrm{ij}}=\frac{\mathrm{C}_{\mathrm{ij}}}{\sigma(\theta_{\mathrm{i}})\sigma(\theta_{\mathrm{j}})}\,.$ (11)
$\mathcal{L}(\bm{\mu}|\bm{X})=\frac{\exp\Bigl{(}-\frac{1}{2}(\bm{X}-\bm{\mu})^{\mathrm{T}}\bm{\Sigma}^{-1}(\bm{X}-\bm{\mu})\Bigl{)}}{\sqrt{(2\pi)^{\mathrm{k}}|\bm{\Sigma}|}}\,,$ (12)
where $\bm{X}$ is the observed $k$-dimensional column vector, $\bm{\mu}$ is the model-predicted $k$-dimensional column vector, $\bm{\Sigma}$ is the symmetric input covariance matrix, and $|\bm{\Sigma}|$ is the determinant of the input covariance matrix.
In the case of 21-cm cosmology, the log-likelihood function is
$\ln\mathcal{L}(\bm{\theta}|\bm{X})\propto-\frac{1}{2}\Bigl{(}(\bm{X}-T(\bm{\theta}))^{T}\bm{\Sigma}^{-1}(\bm{X}-T(\bm{\theta}))\Bigl{)}\,,$ (13)
where $\bm{\theta}$ is the parameter column vector and $T(\bm{\theta})$ is the temperature of the global 21-cm signal generated by ares code for $\bm{\theta}$ as its input parameters. In the following subsection, we will elaborate more on the models used for the input covariance matrix.
III.2 Noise Models
@@
Ref. [69] demonstrated that that only a limited number of frequency channels from the EDGES experiment are effectively independent. The reason is the presence of structures with a period of less than $\sim 10$ MHz in the residual plot of EDGES data after subtracting the foreground and the template of the 21-cm signal. Using the same method employed in Ref. [69] to estimate spectral structures, we identified features with a characteristic scale of less than $\sim 6$ MHz in the SARAS data. Based on this, we assume ten independent data points for both EDGES-like and SARAS-like experiments. Accordingly, we model the input covariance matrix as diagonal, implying no correlations between different frequency channels.
III.2.1 RMS (White) Noise
As mentioned above, the RMS values for the noise level for each experiment depend on the smoothing frequency. Specifically, the RMS values are $20.2$ mK and $23.4$ mK at smoothing frequencies of $3.12$ MHz and $2.8$ MHz for the EDGES [62] and SARAS3 [63] experiments, respectively. These values are consistent with the data presented in Fig. $5$b of Ref. [63]. The higher noise levels in the SARAS3 experiment can primarily be attributed to its lower integration time, though improvements are expected in the future. We use these RMS noises in our analysis for EDGES-like and SARAS-like experimental setups (see the first two rows in Table 1 for details).
III.2.2 Radiometer Noise
The noise of an ideal radiometer with an integration time of $\tau$ and channel width $\Delta\nu$ with a system temperature $\mathrm{T}_{\mathrm{sys}}$ is given by [105, 106]
$\sigma_{\mathrm{T}}=\frac{\mathrm{T}_{\mathrm{sys}}}{\sqrt{\Delta\nu\ \tau}}\,,$ (14)
where $\mathrm{T}_{\mathrm{sys}}=\mathrm{T}_{\mathrm{sky}}+\mathrm{T}_{\mathrm{receiver}}$ with $T_{\mathrm{sky}}$ and $\mathrm{T}_{\mathrm{receiver}}$ being the temperatures of sky and receiver, respectively. We calculate the sky temperature using the $408$ MHz Haslam all-sky map [107], which is defined as
$\mathrm{T}_{\mathrm{sky}}=\mathrm{T}_{408}\left(\frac{\nu}{\nu_{408}}\right)^{\beta}\ ,$ (15)
where we fix $\beta=-2.56$ and $\mathrm{T}_{408}=22.7$ K. Following the calculations presented in Ref. [72], we consider $\mathrm{T}_{\mathrm{receiver}}\approx 300$ K, $\tau=107$ h, $\Delta\nu=390.625$ kHz for the EDGES-like experiment. Since Future1 and Future2 experimental setups are enhanced versions of the EDGES-like experiment, featuring longer integration times and additional frequency channels, we also apply the same specifications to these scenarios.
@@ III.2.2 Radiometer Noise
Fig. 2 shows the input noise as a of function frequency for both noise models. It is evident from this figure that the white noise levels in a SARAS-like at present stage are comparable to that in an EDGES-like experiment. According to Eq. 14, the critical parameter to reduce the value of radiometer noise is the product of the channel width and the integration time. In other words, reducing the channel width requires increasing the integration time to keep a constant noise level, assuming all other parameters remain unchanged. This suggests that a higher integration time is needed for SARAS-like experiments to maintain the same level of radiometer noise compared to an EDGES-like experiment. Hence, for our analysis, we consider only the noise levels associated with EDGES for our futuristic scenarios. However, the pipeline presented here can be easily adapted to accommodate different noise levels for future and ongoing experiments.
@@
$\mathrm{f}_{\mathrm{*}}$ 0.01 (fixed)
$\mathrm{f}_{\mathrm{esc}}$ 0.1
$\mathrm{f}_{\mathrm{X}}$ 1.0
$\mathrm{T}_{\mathrm{min}}$ 500
$\mathrm{N}_{\mathrm{lw}}$ 9690
$\mathrm{m}_{\mathrm{\chi}}$ 100 kev-1 Tev
@@
IV Results
Throughout this paper, we assume $\Lambda$CDM cosmology as our fiducial model. Given the discrepancies between the results reported by the EDGES and SARAS3 experiments, we do not attempt to analyze their data sets. Instead, we use the two experiments as realistic experimental test cases to assess the sensitivity of the global signal to DM physics. Our approach is to generate a mock realization of the global 21-cm signal within the $\Lambda$CDM cosmology and assess the sensitivity of each experimental configuration, assuming noise levels similar to those of EDGES and SARAS3.
We assume that the observed signal is consistent with the standard $\Lambda$CDM cosmology predictions and infer the upper bound on DM-baryon interaction cross section in the case of a null detection. However, the pipeline presented in this work is designed to be flexible and can be easily extended to incorporate different noise and fiducial models. This adaptability ensures that the framework remains applicable for future explorations of the data as noise levels decrease in future observations.
As explained in Section II, we utilize the modified version of three codes, class, galacticus, and ares, as a single pipeline to generate the global 21-cm signals within the IDM cosmology. We then use these generated signals to perform the Fisher forecast analysis with two different input covariance noise models to estimate the sensitivities of four experimental scenarios. An example of such global 21-cm signals is presented in Fig. 1.
@@ IV Results
We perform a multi-parameter Fisher forecasting analysis with the following set of astrophysical and IDM parameters: $\bm{\theta}=[\mathrm{f}_{\mathrm{esc}},\mathrm{f}_{\mathrm{X}},\mathrm{T}_{\mathrm{min}},\mathrm{N}_{\mathrm{lw}},\sigma_{\mathrm{0}}]$. The fiducial values for each parameter are summarized in Table 2, and additional details about these parameters are provided in Section II. It is important to note that we have kept $\mathrm{f}_{\mathrm{\ast}}$ fixed, as it is highly degenerate with nearly all other astrophysical parameters considered in this analysis. After calculating the Fisher matrix, we use fishchips  [^4] https://github.com/xzackli/fishchips-public [108] to visualize the 2-D and 1-D marginalized posteriors. These results for IDM models with $n=0$ and $n=-4$ are shown in Figs. 3 and 4, respectively.
@@
Figure 6: The 95% confidence level upper limit forecasts on $\log_{10}(\sigma_{0}/\mathrm{cm}^{2})$ as a function of IDM mass are shown for Coulomb-like IDM model ($n=-4$) and four different experimental scenarios (the shaded region, or area above a given sensitivity curve, would be excluded by a non-detection of a signal with a given experiment). The previous bounds from CMB [28] are also plotted for comparison. All four experimental scenarios predict stronger forecasts compared to the current CMB bounds.
@@
Fig. 3 presents the two-dimensional $68\%$, $95\%$ confidence contours for all parameters of interest, showing the correlation between different astrophysical parameters and the interaction cross section for the $n=0$ IDM model. The top panel in each column of the figure shows the marginalized posterior distributions for the corresponding parameters. It is evident from the plot that $\sigma_{\mathrm{0}}$ has no correlation with $\mathrm{f}_{\mathrm{X}}$ and features an anti-correlation with $\mathrm{T}_{\mathrm{min}}$. On the one hand, this suggests that the interaction cross section can be reconstructed with little bias from astrophysical uncertainty related to the X-ray luminosity of early galaxies. On the other hand, the increase in the minimum virial temperature delays the formation of the first galaxies, causing a delay in the heating and ionization of the intergalactic medium, which shifts the trough of the 21-cm signal to higher frequencies; a velocity-independent interaction between DM and baryons affects the 21-cm signal similarly. A minimum virial temperature, related to the minimum mass of halos forming galaxies, is therefore a major source of uncertainty in DM interaction inference with the global 21-cm signal. We further find a significant positive correlation between $\sigma_{\mathrm{0}}$ and $\mathrm{N}_{\mathrm{lw}}$, suggesting that an increase in the interaction cross section requires a higher number of Lyman-Werner photons to compensate for the signal shift caused by the enhanced interaction rate. The uncertainty in $\mathrm{N}_{\mathrm{lw}}$ is therefore likewise of concern to DM inference with the global signal. Finally, the broad uncertainties in the forecasts for $\mathrm{f}_{\mathrm{esc}}$ indicate that the global 21-cm signal is largely insensitive to this parameter within the frequency range considered across the four scenarios. This is because $\mathrm{f}_{\mathrm{esc}}$ primarily affects the reionization process, which is expected to occur mostly at $\mathrm{z}\lesssim 10$ based on other observational constraints, outside the most sensitive redshift range probed here.
@@
Fig. 4 presents the two-dimensional $68\%$, $95\%$ confidence contours, showing the forecasts for different astrophysical parameters and the cross section of interaction for the IDM model with Coulomb-like interaction ($n=-4$). We find a positive correlation between $\sigma_{\mathrm{0}}$ and $\mathrm{f}_{\mathrm{X}}$, as both parameters tend to reduce the depth of the global signal and move it to lower frequencies (for low enough cross sections). Furthermore, increasing the cross section delays the structure formation, which is degenerate with a decrease in the number of Lyman-Werner photons ($\mathrm{N}_{\mathrm{lw}}$). However, if baryons cool through interactions with DM, a lower virial temperature, an increase in the number of Lyman-Werner photons, and an increase in X-ray efficiency is needed to preserve the amplitude of the signal. We also note a strong anti-correlation between $\mathrm{T}_{\mathrm{min}}$ and $\mathrm{N}_{\mathrm{lw}}$, indicating that the two parameters are difficult to constrain independently with global signal data alone. Similar to the $n=0$ case, $\mathrm{f}_{\mathrm{esc}}$ is not highly correlated with any other parameters, as its impact on the signal becomes significant only at high frequencies.
@@
Fig. 5 is derived by repeating the analysis represented by Fig. 3, for a range of DM particle masses, and marginalizing over all astrophysical parameters. This figure shows a forecast of the $95\%$ confidence level upper bound on the interaction cross section between DM and baryons for the $n=0$ model, as a function of the DM particle mass. Forecasts are computed for four different experimental scenarios, as listed in Table 1 and discussed in Subsection III.2. The region of the parameter space above each curve would be excluded at the $95\%$ confidence level by a given experiment. For example, the shaded region in the plot corresponds to the region excluded by the EDGES-like configuration. The current bounds from CMB Planck data (dotted grey) [28] and from the Milky Way satellite population measurements (dot-dashed grey) [23] are also shown for comparison. The light blue and dark blue curves correspond to constraints from SARAS-like and EDGES-like scenarios, respectively. Notably, the SARAS-like scenario, which has the lowest sensitivity among the four cases explored in this work, can achieve a comparable upper bound to the best existing bounds derived using the Milky Way satellite data [23]. The light and dark dotted green curves show the forecasts for Future1 and Future2 experimental setups, respectively. These forecasts illustrate the improvement in sensitivity with increasing observation time and more frequency channels.
@@
Fig. 6 similarly shows the 95% forecasts on the cross section of interaction between DM and baryons for the $n=-4$ model, for a range of DM particle masses. Similar to the forecast plot for the $n=0$ model, these forecasts are computed for four different experimental scenarios. The previous bounds from CMB (dotted grey) [28] are also plotted for comparison. The light and dark blue curves correspond to the constraints for SARAS-like and EDGES-like scenarios. To emphasize the forecasts for the EDGES-like scenario, the excluded parameter space is shaded in blue. It can also be seen in Fig. 6 that the global signal is highly sensitive to the $n=-4$ model (see also Fig. 1). Consequently, the SARAS-like experiment can in principle surpass the sensitivity of the CMB data to these interaction scenarios. This improvement is particularly pronounced for lower masses, where the forecasted upper bound on the interaction cross section is more than an order of magnitude lower than the current observational bounds. Additionally, we show the forecasts for two futuristic scenarios in the light-dotted and dark-dotted green curves to highlight the potential of future 21-cm measurements in exploring DM interactions. The corresponding projected 95% confidence level upper limits are also presented in Table 3.
V Conclusions

### v2

@@ 3.1 Fisher Forecasting
Table 1: Comparison of four experimental setups, outlining their key parameters such as minimum and maximum frequencies probed ($\nu_{\mathrm{min}},\nu_{\mathrm{max}}$), number of independent frequency channels ($N_{\mathrm{ch}}$), native (not smoothed) frequency bins ($\Delta\nu$), observation times ($\tau_{\mathrm{obs}}$), receiver temperatures ($T_{\mathrm{rec}}$), and RMS value of the smoothed noise ($\sigma_{\mathrm{RMS}}$).
3.2 Noise Models
@@
Ref. [55] demonstrated that only a limited number of frequency channels from the EDGES experiment are effectively independent. The reason is the presence of structures with a period of less than $\sim 10$ MHz in the residual plot of EDGES data after subtracting the foreground and the template of the 21-cm signal. Using the same method employed in Ref. [55] to estimate spectral structures, we identified features with a characteristic scale of less than $\sim 6$ MHz in the SARAS data. Based on this, we assume ten independent data points for both EDGES-like and SARAS-like experiments. Accordingly, we model the input covariance matrix as diagonal, implying no correlations between different frequency channels.
3.2.1 RMS (White) Noise
As mentioned above, the RMS values for the noise level for each experiment depend on the smoothing frequency. Specifically, the RMS values are $20.2$ mK and $23.4$ mK at smoothing frequencies of $3.12$ MHz and $2.8$ MHz for the EDGES [24] and SARAS3 [93] experiments, respectively. These values are consistent with the data presented in Fig. 55b of Ref. [93]. The higher noise levels in the SARAS3 experiment can primarily be attributed to its lower integration time, though improvements are expected in the future. We use these RMS noises in our analysis for EDGES-like and SARAS-like experimental setups (see the first two rows in Table 1 for details).
3.2.2 Radiometer Noise
The noise of an ideal radiometer with an integration time of $\tau$ and channel width $\Delta\nu$ with a system temperature $T_{\mathrm{sys}}$ is given by [103, 47]
$\sigma_{\mathrm{T}}=\frac{T_{\mathrm{sys}}}{\sqrt{\Delta\nu\ \tau}}\,$, (3.7)
where $T_{\mathrm{sys}}=T_{\mathrm{sky}}+T_{\mathrm{receiver}}$ with $T_{\mathrm{sky}}$ and $T_{\mathrm{receiver}}$ being the temperatures of sky and receiver, respectively. We calculate the sky temperature using the $408$ MHz Haslam all-sky map [51], which is defined as
$T_{\mathrm{sky}}=T_{408}\left(\frac{\nu}{\nu_{408}}\right)^{\beta}\ ,$ (3.8)
where we fix $\beta=-2.56$ and $T_{408}=22.7$ K. Following the calculations presented in Ref. [92], we consider $T_{\mathrm{receiver}}\approx 300$ K, $\tau=107$ h, $\Delta\nu=390.625$ kHz for the EDGES-like experiment. Since Future1 and Future2 experimental setups are enhanced versions of the EDGES-like experiment, featuring longer integration times and additional frequency channels, we also apply the same specifications to these scenarios.
@@ 3.2.2 Radiometer Noise
Fig. 2 shows the input noise as a of function frequency for both noise models. It is evident from this figure that the white noise levels in a SARAS-like at present stage are comparable to that in an EDGES-like experiment. According to Eq. 3.7, the critical parameter to reduce the value of radiometer noise is the product of the channel width and the integration time. In other words, reducing the channel width requires increasing the integration time to keep a constant noise level, assuming all other parameters remain unchanged. This suggests that a higher integration time is needed for SARAS-like experiments to maintain the same level of radiometer noise compared to an EDGES-like experiment. Hence, for our analysis, we consider only the noise levels associated with EDGES for our futuristic scenarios. However, the pipeline presented here can be easily adapted to accommodate different noise levels for future and ongoing experiments.
@@
$f_{\mathrm{*}}$ 0.01 (fixed)
$f_{\mathrm{esc}}$ 0.1
$f_{\mathrm{X}}$ 1.0
$T_{\mathrm{min}}$ 500
$N_{\mathrm{lw}}$ 9690
$m_{\mathrm{\chi}}$ 100 kev-1 Tev
@@
4 Results
[…]
As explained in Section 2, we utilize the modified version of three codes, class, galacticus, and ares, as a single pipeline to generate the global 21-cm signals within the IDM cosmology. We then use these generated signals to perform the Fisher forecast analysis with two different input covariance noise models to estimate the sensitivities of four experimental scenarios. An example of such global 21-cm signals is presented in Fig. 1.
We perform a multi-parameter Fisher forecasting analysis with the following set of astrophysical and IDM parameters: $\boldsymbol{\theta}=[f_{\mathrm{esc}},f_{\mathrm{X}},T_{\mathrm{min}},N_{\mathrm{lw}},\sigma_{\mathrm{0}}]$. The fiducial values for each parameter are summarized in Table 2, and additional details about these parameters are provided in Section 2. It is important to note that we have kept $f_{\mathrm{\ast}}$ fixed, as it is highly degenerate with nearly all other astrophysical parameters considered in this analysis. After calculating the Fisher matrix, we use fishchips  [^4] https://github.com/xzackli/fishchips-public [59] to visualize the 2-D and 1-D marginalized posteriors. These results for IDM models with $n=0$ and n=4n=-4 are shown in Figs. 3 and 4, respectively.
@@ 4 Results
Figure 6: The 95% confidence level upper limit forecasts on $\log_{10}(\sigma_{0}/\mathrm{cm}^{2})$ as a function of IDM mass are shown for Coulomb-like IDM model (n=4n=-4) and four different experimental scenarios (the shaded region, or area above a given sensitivity curve, would be excluded by a non-detection of a signal with a given experiment). The previous bounds from CMB [76] are also plotted for comparison. All four experimental scenarios predict stronger forecasts compared to the current CMB bounds.
@@
Fig. 3 presents the two-dimensional $68\%$, $95\%$ confidence contours for all parameters of interest, showing the correlation between different astrophysical parameters and the interaction cross section for the $n=0$ IDM model. The top panel in each column of the figure shows the marginalized posterior distributions for the corresponding parameters. It is evident from the plot that $\sigma_{\mathrm{0}}$ has no correlation with $f_{\mathrm{X}}$ and features an anti-correlation with $T_{\mathrm{min}}$. On the one hand, this suggests that the interaction cross section can be reconstructed with little bias from astrophysical uncertainty related to the X-ray luminosity of early galaxies. On the other hand, the increase in the minimum virial temperature delays the formation of the first galaxies, causing a delay in the heating and ionization of the intergalactic medium, which shifts the trough of the 21-cm signal to higher frequencies; a velocity-independent interaction between DM and baryons affects the 21-cm signal similarly. A minimum virial temperature, related to the minimum mass of halos forming galaxies, is therefore a major source of uncertainty in DM interaction inference with the global 21-cm signal. We further find a significant positive correlation between $\sigma_{\mathrm{0}}$ and $N_{\mathrm{lw}}$, suggesting that an increase in the interaction cross section requires a higher number of Lyman-Werner photons to compensate for the signal shift caused by the enhanced interaction rate. The uncertainty in $N_{\mathrm{lw}}$ is therefore likewise of concern to DM inference with the global signal. Finally, the broad uncertainties in the forecasts for $f_{\mathrm{esc}}$ indicate that the global 21-cm signal is largely insensitive to this parameter within the frequency range considered across the four scenarios. This is because $f_{\mathrm{esc}}$ primarily affects the reionization process, which is expected to occur mostly at $z\lesssim 10$ based on other observational constraints, outside the most sensitive redshift range probed here.
@@
Fig. 4 presents the two-dimensional $68\%$, $95\%$ confidence contours, showing the forecasts for different astrophysical parameters and the cross section of interaction for the IDM model with Coulomb-like interaction (n=4n=-4). We find a positive correlation between $\sigma_{\mathrm{0}}$ and $f_{\mathrm{X}}$, as both parameters tend to reduce the depth of the global signal and move it to lower frequencies (for low enough cross sections). Furthermore, increasing the cross section delays the structure formation, which is degenerate with a decrease in the number of Lyman-Werner photons ($N_{\mathrm{lw}}$). However, if baryons cool through interactions with DM, a lower virial temperature, an increase in the number of Lyman-Werner photons, and an increase in X-ray efficiency is needed to preserve the amplitude of the signal. We also note a strong anti-correlation between $T_{\mathrm{min}}$ and $N_{\mathrm{lw}}$, indicating that the two parameters are difficult to constrain independently with global signal data alone. Similar to the $n=0$ case, $f_{\mathrm{esc}}$ is not highly correlated with any other parameters, as its impact on the signal becomes significant only at high frequencies.
@@
Fig. 5 is derived by repeating the analysis represented by Fig. 3, for a range of DM particle masses, and marginalizing over all astrophysical parameters. This figure shows a forecast of the $95\%$ confidence level upper bound on the interaction cross section between DM and baryons for the $n=0$ model, as a function of the DM particle mass. Forecasts are computed for four different experimental scenarios, as listed in Table 1 and discussed in Subsection 3.2. The region of the parameter space above each curve would be excluded at the $95\%$ confidence level by a given experiment. For example, the shaded region in the plot corresponds to the region excluded by the EDGES-like configuration. The current bounds from CMB Planck data (dotted grey) [76] and from the Milky Way satellite population measurements (dot-dashed grey) [63] are also shown for comparison. The light blue and dark blue curves correspond to constraints from SARAS-like and EDGES-like scenarios, respectively. Notably, the SARAS-like scenario, which has the lowest sensitivity among the four cases explored in this work, can achieve a comparable upper bound to the best existing bounds derived using the Milky Way satellite data [63]. The light and dark dotted green curves show the forecasts for Future1 and Future2 experimental setups, respectively. These forecasts illustrate the improvement in sensitivity with increasing observation time and more frequency channels.
@@
Fig. 6 similarly shows the 95% forecasts on the cross section of interaction between DM and baryons for the n=4n=-4 model, for a range of DM particle masses. Similar to the forecast plot for the $n=0$ model, these forecasts are computed for four different experimental scenarios. The previous bounds from CMB (dotted grey) [76] are also plotted for comparison. The light and dark blue curves correspond to the constraints for SARAS-like and EDGES-like scenarios. To emphasize the forecasts for the EDGES-like scenario, the excluded parameter space is shaded in blue. It can also be seen in Fig. 6 that the global signal is highly sensitive to the n=4n=-4 model (see also Fig. 1). Consequently, the SARAS-like experiment can in principle surpass the sensitivity of the CMB data to these interaction scenarios. This improvement is particularly pronounced for lower masses, where the forecasted upper bound on the interaction cross section is more than an order of magnitude lower than the current observational bounds. Additionally, we show the forecasts for two futuristic scenarios in the light-dotted and dark-dotted green curves to highlight the potential of future 21-cm measurements in exploring DM interactions. The corresponding projected 95% confidence level upper limits are also presented in Table 3. We can also translate the constraints on $\sigma_{0}$ to the physical cross-section by multiplying with the factor of $\left(v_{\mathrm{rel}}/c\right)^{n}$, where $c$ is the speed of light. For $n=0$, $\sigma_{0}$ is identical to the physical cross section, while for Coulomb-like interactions they differ by a factor of $\left(v_{\mathrm{rel}}/c\right)^{-4}$. For completeness, we provide the multiplicative factors required to convert $\sigma_{0}$ to the physical cross section $\sigma$ at two representative epochs: $z=0$ and $z=18$. The latter corresponds to the cosmic epoch near the absorption trough of the global 21-cm signal reported by the EDGES collaboration, while $z=0$ represents the present day for reference. At redshifts $z=0$ and $z=18$, the relative bulk velocities are $v_{\mathrm{rel}}=29.97~\mathrm{m/s}$ and $v_{\mathrm{rel}}=569.4~\mathrm{m/s}$, respectively, yielding multiplicative conversion factors of $1.00\times 10^{28}$ and $7.68\times 10^{22}$, respectively. The projected 95% confidence level upper limits on the physical cross section for each experimental configuration can thus be obtained by applying these corresponding factors to the reported constraints on $\sigma_{0}$.
5 Conclusions